Two monkeys of masses 10 kg and 8 kg are moving along a vertical light rope, the former climbing up with an acceleration of 2m/s 2 while the latter coming down with a uniform velocity of 2 m/s. Find the tension in the rope at the fixed support. need the procedure ..

Grade:12

2 Answers

sahastra kishore
7 years ago
Tension due to monkey on A =m (g+a) 10 (10+2) =120Tension due to monkey B =mg =8×10 =80Total tension =(120+80) =200N
Vardaan
2 years ago

Tension due to monkey on A=m(g+a)

=10(10+2)=120
Tension due to monkey B=mg=8×10=80
Total tension =(120+80)=200N